# Reconfigurable dissipative entanglement between many spin ensembles: from robust quantum sensing to many-body state engineering

### v1

@@ Abstract
An attractive approach for stabilizing entangled many-body spin states is to employ engineered dissipation. Most existing proposals either target relatively simple collective spin states, or require numerous independent and complex dissipative processes. Here, we show a surprisingly versatile scheme for many-body reservoir engineering that relies solely on fully collective single-excitation decay, augmented with local Hamiltonian terms. Crucially, all these ingredients are readily available in cavity QED setups. Our method is based on splitting the spin system into groups of sub-ensembles, and provides an easily tunable setup for stabilizing a broad family of pure, highly entangled states with closed-form analytic descriptions. Our results have immediate application to multi-ensemble quantum metrology, enabling Heisenberg-limited sensing of field gradients and curvatures. Notably, the generated states have robustness against common-mode phase noise, and only require simple Ramsey-style measurements. The same setup also allows the stabilization of entangled states in a 1D chain of spin ensembles with symmetry-protected topological (SPT) order, and have a direct connection to the outputs of sequential unitary circuits. In particular, we present an efficient method for engineering the celebrated spin-1 Affleck-Kennedy-Lieb-Tasaki (AKLT) state.
@@ I Introduction
Cavity QED platforms also naturally provide an alternative route for generating nontrivial spin dynamics and entangled states by harnessing cavity dissipation. A generic situation in experiments is the limit where cavity loss induces fully collective dissipation on the spins, a single dissipative process where each atom is essentially indistinguishable. This setting underlies the celebrated phenomenon of Dicke superradiance [3, 4], characterized by a collectively enhanced atomic decay and the generation of a photonic burst. Beyond superradiance, the interplay between collective decay and coherent Rabi driving can give rise to dissipative phase transitions [5, 6] and even time-crystal–like phenomena [7]. Moreover, the combination of collective decay and tailored external driving enables the generation of metrologically useful spin-squeezed states [8, 9, 10], as well as extensions to two-ensemble configurations [11, 12, 13, 14].
@@
While permutation-symmetric spin models with collective cavity-induced decay are appealing for their experimental relevance, a broader goal is to realize more general forms of reservoir engineering [15], where controlled dissipation can stabilize a wide variety of nontrivial many-body spin states beyond the permutation-symmetric case. Many-body dissipative state-preparation schemes are of interest both for their practical utility, such as the potential to design protocols that are more robust than unitary approaches, and for their fundamental implications including the characterization of dissipative phases of matter [16, 17]. There are by now a host of theoretically proposed protocols for reservoir-engineering (see, e.g., Ref. 18 for a recent review). However, these protocols typically involve a considerable amount of complexity for near-term experiments: Instead of a single collective cavity-induced loss process, they often require engineering a large, even extensive, number of independent and highly tailored dissipative processes.
@@
Given this context, a natural question is whether one can bridge these two extremes: is it possible to only rely on the ubiquitous ingredient of collective loss in cavity QED systems, but still dissipatively stabilize complex, non-collective many-body states? One might even be more ambitious and ask whether such an approach could be made reconfigurable, allowing the same setup to stabilize a wide variety of target states. In this work, surprisingly we show that this is indeed possible. We introduce and analyze a general method for dissipative stabilization of a broad class of non-collective many-body entangled states using only a single collective loss dissipator. Our approach breaks the permutation symmetry of $N_{\rm tot}$ atoms only through Hamiltonian terms, namely a pattern of drive detunings and a structured cavity-mediated spin-exchange interaction. The Hamiltonian terms reduce the symmetry of the system and lead to $L$ distinct spin ensembles, each with the same collective spin $S=N_{\rm tot}/2L$. By varying the detuning pattern, we show that one can dissipatively stabilize a wide range of pure entangled states of these $L$ ensembles. We can further derive closed-form analytic descriptions of the corresponding steady states. Note that in the extreme limit $L=N_{\rm tot}$, where the steady states attain a greatly simplified structure, our construction coincides with Refs. 19, 20, 21. Our work has immediate relevance to recent cavity QED experiments involving two or more spin ensembles [22, 23, 24, 25].
@@
While the space of states accessible with our approach is vast, we focus on two particularly intriguing classes. In Secs. III and IV, we demonstrate how multi-ensemble entangled states with exceptional metrological properties can be naturally stabilized. These states enable Heisenberg-limited sensing of multiple parameters such as field gradients and curvatures, exhibit near-complete insensitivity to common-mode noise, and allow the full quantum Fisher information (QFI) to be extracted using a simple Ramsey-like protocol. In Sec. V, we discuss how our approach can stabilize states of fundamental interest in many-body physics: a family of entangled states in a 1D chain of spin ensembles that can exhibit symmetry-protected topological (SPT) order [26, 27], and that connect directly to recent ideas for exploiting sequential unitary circuits [28]. In particular, we present an efficient method for stabilizing the celebrated spin-1 Affleck–Kennedy–Lieb–Tasaki (AKLT) state [29, 30]. Together, these examples highlight both the power and versatility of our general approach.
@@ II Setup for reconfigurable reservoir engineering
Our approach is sketched in Fig. 1. We start with a simple, fully-collective system where $N_{\rm tot}$ two-level atoms experience a single collective decay process (induced by resonant interaction with a lossy cavity), and are also subject to uniform and resonant Rabi drives. This realizes the well-known cooperative Rabi fluorescence (CRF) model [5, 6]. Working in the rotating frame and adiabatically eliminating the cavity, the spin dynamics follows a Gorini–Kossakowski–Sudarshan–Lindblad (Lindblad) master equation:
@@
The CRF model is fully collective (i.e., total angular momentum is conserved), severely limiting the complexity of the resulting dynamics and steady states. As promised, we will achieve a richer dynamics by perturbing the system Hamiltonian, while keeping the dissipative part of the dynamics unchanged (see Fig. 1). We first partition our full set of $N_{\rm tot}$ atoms equally into $L$ sub-ensembles (indexed by $l=1,\dots,L$, with $L$ an even number). Then we modify the Rabi drives so that atoms in sub-ensemble $l$ experience a drive-detuning $\delta_{l}$. We also introduce an all-to-all spin-exchange interaction (amplitude $\chi$) with a “chiral” structure [19, 20, 21]. The modified master equation describes a system of $L$ ensembles with the same collective spin $S=N_{\rm tot}/2L$, and has the form:
@@
All the new terms introduced in $\hat{H}$ are compatible with the experimental tools of cavity QED. As shown in Fig. 1, one can implement Eq. (2) by trapping many spin ensembles in a cavity, each with different detunings $\delta_{l}$. Similar to the setup in Ref. 22, the chiral spin-exchange couplings $\chi$ can be engineered by applying a magnetic-field gradient along the cavity, which allows for engineering cavity-assisted Raman processes between ensembles separated by different distances independently via frequency selection. Alternatively, one can also engineer $\chi$ using a chiral quantum network [19, 20, 21, 36, 37], which functions by coupling each spin ensemble to an optical cavity, and then connecting all these cavities via chiral waveguides. We discuss the details of the implementation of chiral spin-exchange couplings $\chi$ in App. G.
@@
Note that our approach is connected to (but distinct from) previous work exploiting non-reciprocal interactions for entanglement stabilization. In the limit where the chiral interaction is tuned to exactly $\chi=\Gamma$, Eq. (2) has the form of a cascaded master equation [38, 39, 19], and mimics a setup where the $L$ ensembles are all coupled to a common unidirectional waveguide (meaning that ensemble $l$ is only influenced by ensembles $k<l$). In the case $L=2$, dissipative pure-state entanglement in this fully directional limit can be understood as an example of the coherent quantum absorber method [19] and a consequence of hidden time-reversal symmetry [40]; applications to two spin ensembles was studied in Ref. 33, 41. Our work shows that for $L=2$, entanglement generation survives beyond the perfectly directional limit, indicating a kind of adiabatic continuity of the steady state as one tunes $\chi$.
In the other extreme case $L=N_{\rm tot}$ (i.e., each ensemble is just a single qubit with $S=1/2$), our work reduces to the entanglement stabilization scheme of Refs. 19, 20, 21. Note that, in this case, the complexity of the stabilized states is reduced, and there is no obvious metrological utility to the states that are generated. Although the possibility of states with SPT order exists for the case of $S=1/2$, it was not discussed by the previous works. Our work non-trivially generalizes this scheme to the case where each subsystem has an arbitrary spin $S>1/2$.
@@ III.2 Quantum enhancement for differential sensing
Figure 3: (a) Generalized Bloch sphere representation of the $L=2$ steady state. It can be interpreted as a two-mode spin squeezed state with squeezing along the $\hat{S}^{y}_{1}+\hat{S}^{y}_{2}$ and $\hat{S}^{z}_{1}+\hat{S}^{z}_{2}$ axes, and anti-squeezing along the $\hat{S}^{z}_{1}-\hat{S}^{z}_{2}$ and $\hat{S}^{y}_{1}-\hat{S}^{y}_{2}$ axes. Applying a differential phase $\phi$ between the two ensembles is equivalent to rotations about the $\hat{S}^{z}_{1}-\hat{S}^{z}_{2}$ axis in the first Bloch sphere. (b) Comparison between the optimal time scale to reach the steady state ($t_{\rm ss}$) and the steady-state squeezing (described by $\mathrm{Var}(\hat{S}^{y}_{1}+\hat{S}^{y}_{2})_{\rm ss}$), with $S=30$. We consider the initial state where all the spins are pointing down, and define $t_{\rm ss}$ as the optimal evolution time for the infidelity to the steady state to reach $10^{-3}$ with fixed $\Gamma$ and $\Delta/\Omega$. The same initial state is used in (c). (c) The scaling of optimal two-mode Wineland spin-squeezing parameter $\xi^{2}_{\rm opt}$ with spin $S$ for each ensemble and single-atom cooperativity $C$ (obtained from a second-order cumulant expansion). The scaling $\xi^{2}_{\rm opt}\propto 1/\sqrt{SC}$ is determined via numerical fitting.
@@
A central question is how this QFI scales with the number of atoms, i.e., with $S$. Fig. 2(d) shows this scaling for different choices of the parameter $\Delta/\Omega$. As discussed in the previous section, away from the large drive regime, i.e., when $\Delta/\Omega\gg 1/S$, the steady state’s angular-momentum distribution is localized near $J\sim S\Delta/\Omega$. It then follows from Eq. (10) and the form of the $B_{J,-J}$ coefficients that $\mathcal{F}_{\phi}\sim S\Omega/\Delta$. This immediately suggests how to achieve different canonical scalings of the QFI. For $\Delta/\Omega\sim 1$, we have a scaling near the standard quantum limit (SQL), $\mathcal{F}_{\phi}\sim S$. For a Rabi drive that increases with system size, $\Delta/\Omega\sim 1/\sqrt{S}$, we obtain a scaling that exceeds the SQL, $\mathcal{F}_{\phi}\sim S^{3/2}$.
For even larger Rabi drives, $\Delta/\Omega\sim 1/S$, we obtain a QFI that exhibits Heisenberg-limited (HL) scaling, i.e., $\mathcal{F}_{\phi}\propto S^{2}$. This is consistent with our earlier observation that our state approaches the $|J=0,m=0\rangle$ state in the large-Rabi-drive limit ($\Omega\rightarrow\infty$), a state which has a maximal QFI $\mathcal{F}_{\phi}]_{\rm max}=16S(S+1)/3$. The same maximal QFI is also reported in Refs. 13, 14. Since the $|J=0,m=0\rangle$ state is invariant under global rotations, in the large-Rabi-drive limit, one can obtain the same HL-scaling QFI for parameter estimation corresponding to generators $\hat{S}^{x}_{1}-\hat{S}^{x}_{2}$, $\hat{S}^{y}_{1}-\hat{S}^{y}_{2}$, and $\hat{S}^{z}_{1}-\hat{S}^{z}_{2}$, and vanishing variances for the operators $\hat{S}^{x}_{1}+\hat{S}^{x}_{2}$, $\hat{S}^{y}_{1}+\hat{S}^{y}_{2}$, and $\hat{S}^{z}_{1}+\hat{S}^{z}_{2}$. The squeezing and anti-squeezing of these operators can be extended to the case of a finite $\Omega$ as we show later, although the rotation symmetry of the three directions no longer holds.
@@
where $B_{J,m}^{-}=\langle J+1,m-1|\hat{S}^{-}_{1}-\hat{S}^{-}_{2}|J,m\rangle$. Since $\mathcal{F}_{\phi}\sim S\Omega/\Delta$, as discussed previously, we have $\langle\hat{S}^{x}_{1}-\hat{S}^{x}_{2}\rangle_{\rm ss}\propto S$ for a wide range of $\Delta/\Omega$, with corrections when decreasing $\Delta/\Omega$ below $\sim 1/S$. This result suggests the possibility of having the QFI exhibit Heisenberg-scaling while still having a non-zero average spin length of order $S$, e.g., by having $\Delta/\Omega\sim 1/S$. We indeed find that this is true through an explicit evaluation of Eq. (11). Note that having a spin length of order $S$ not only ensures that a Ramsey-type protocol is applicable, but also enhances the robustness against detection noise via large measurable signals.
Based on the large spin length $\langle\hat{S}^{x}_{1}-\hat{S}^{x}_{2}\rangle_{\rm ss}\propto S$ in our case, as shown in Fig. 3(a), one can further interpret our steady state $|\psi_{\rm ss}^{(2)}\rangle$ as a two-mode spin squeezed state (see Refs. 43, 44, 13 for other types of two-mode spin squeezed states). The squeezed quadratures are along the $\hat{S}^{y}_{1}+\hat{S}^{y}_{2}$ and $\hat{S}^{z}_{1}+\hat{S}^{z}_{2}$ axes, and the anti-squeezed quadratures are along the $\hat{S}^{z}_{1}-\hat{S}^{z}_{2}$ and $\hat{S}^{y}_{1}-\hat{S}^{y}_{2}$ axes. Both of the Bloch spheres share the same spin length, which is along the $\hat{S}^{x}_{1}-\hat{S}^{x}_{2}$ axis. In particular, the differential phase $\phi$ leads to rotations along the $\hat{S}^{z}_{1}-\hat{S}^{z}_{2}$ axis, resulting in a non-zero average value of $\hat{S}^{y}_{1}+\hat{S}^{y}_{2}$. This suggests a simple Ramsey-style measurement to measure the average collective spin component $\hat{S}^{y}_{1}+\hat{S}^{y}_{2}$. The quantum enhancement of this Ramsey-type measurement can be characterized by the two-mode generalization of the Wineland spin squeezing parameter [13]. Letting $(\Delta\phi)_{\rm est}$ denote the error in estimating $\phi$ from the Ramsey measurement, we have
@@
Note that $\mathrm{Var}(\hat{S}^{y}_{1}+\hat{S}^{y}_{2})_{\rm ss}=\sum_{J=0}^{2S}|c_{J,-J}|^{2}J/2$. Based on the previous discussions of $|c_{J,-J}|^{2}$, one can obtain $\xi^{2}\sim\Delta/\Omega$ when $\Delta/\Omega\gg 1/S$. Thus, by setting $\Delta/\Omega\sim 1/S$, one finds that Eq. (12) is approaching HL scaling, $\xi_{\rm ss}^{2}\sim 1/S$. We thus have established a major advantage of our scheme: HL scaling can be achieved for the differential phase $\phi$ using a simple Ramsey-type measurement with a signal of size $O(S)$. Based on the Bloch sphere representation, one can also extract the phase imprinted by $\hat{S}^{y}_{1}-\hat{S}^{y}_{2}$ with HL scaling by measuring $\hat{S}^{z}_{1}+\hat{S}^{z}_{2}$.
Our approach and the resulting steady state offer several distinct advantages over previous dissipative preparation schemes for enhanced differential sensing based on entanglement between two spin ensembles. Ref. 14 stabilizes a state with vanishing spin length, precluding its use in Ramsey-type measurements. Ref. 13 stabilizes a similar (but distinct) two-mode squeezed state, while it relies on engineering two independent dissipators, making the setup more demanding experimentally. In contrast, our scheme is naturally compatible with a simple Ramsey-type protocol and achieves HL scaling with a simpler experimental design.

### v2

@@ Abstract
An attractive approach for stabilizing entangled many-body spin states is to employ engineered dissipation. Most existing proposals either target relatively simple collective spin states, or require numerous independent and complex dissipative processes. Here, we show a surprisingly versatile scheme for many-body reservoir engineering that relies solely on fully collective single-excitation decay, augmented with local Hamiltonian terms. Crucially, all these ingredients are readily available in cavity QED setups. Our method is based on splitting the spin system into groups of sub-ensembles, and provides an easily tunable setup for stabilizing a broad family of pure, highly entangled states with closed-form analytic descriptions. Our results have immediate application to multi-ensemble quantum metrology, enabling Heisenberg-limited sensing of field gradients and curvatures. Notably, our approach solves an important challenge in differential quantum sensing by providing the first example of Heisenberg-limited differential sensing immune to common-mode noise and accessible with only simple one-body measurements. The same setup also allows the stabilization of an entire family of entangled states in a 1D chain of spin ensembles with symmetry-protected topological (SPT) order, and have a direct connection to the outputs of sequential unitary circuits. A special case of our protocol efficiently stabilizes the celebrated Affleck–Kennedy–Lieb–Tasaki (AKLT) state.
@@ I Introduction
Cavity QED platforms also naturally provide an alternative route for generating nontrivial spin dynamics and entangled states by harnessing cavity dissipation. A generic situation in experiments is the limit where cavity loss induces fully collective dissipation on the spins, a single dissipative process where each atom is essentially indistinguishable. This setting underlies the celebrated phenomenon of Dicke superradiance [3, 4], characterized by a collectively enhanced atomic decay and the generation of a photonic burst. Beyond superradiance, the interplay between collective decay and coherent Rabi driving can give rise to dissipative phase transitions [5, 6] and even time-crystal–like phenomena [7]. Moreover, the combination of collective decay and tailored external driving enables the generation of metrologically useful spin-squeezed states [8, 9, 10, 11, 12], as well as extensions to two-ensemble configurations [13, 14, 15, 16].
@@
While permutation-symmetric spin models with collective cavity-induced decay are appealing for their experimental relevance, a broader goal is to realize more general forms of reservoir engineering [17], where controlled dissipation can stabilize a wide variety of nontrivial many-body spin states beyond the permutation-symmetric case. Many-body dissipative state-preparation schemes are of interest both for their practical utility, such as the potential to design protocols that are more robust than unitary approaches, and for their fundamental implications including the characterization of dissipative phases of matter [18, 19]. There are by now a host of theoretically proposed protocols for reservoir-engineering (see, e.g., Ref. 20 for a recent review). However, these protocols typically involve a considerable amount of complexity for near-term experiments: Instead of a single collective cavity-induced loss process, they often require engineering a large, even extensive, number of independent and highly tailored dissipative processes.
@@
Given this context, a natural question is whether one can bridge these two extremes: is it possible to only rely on the ubiquitous ingredient of collective loss in cavity QED systems, but still dissipatively stabilize complex, non-collective many-body states? One might even be more ambitious and ask whether such an approach could be made reconfigurable, allowing the same setup to stabilize a wide variety of target states. In this work, surprisingly we show that this is indeed possible. We introduce and analyze a general method for dissipative stabilization of a broad class of non-collective many-body entangled states using only a single collective loss dissipator. Our approach breaks the permutation symmetry of $N_{\rm tot}$ atoms only through Hamiltonian terms, namely a pattern of drive detunings and a structured cavity-mediated spin-exchange interaction. The Hamiltonian terms reduce the symmetry of the system and lead to $L$ distinct spin ensembles, each with the same collective spin $S=N_{\rm tot}/2L$. By varying the detuning pattern, we show that one can dissipatively stabilize a wide range of pure entangled states of these $L$ ensembles. We can further derive closed-form analytic descriptions of the corresponding steady states. Note that in the extreme limit $L=N_{\rm tot}$, where the steady states attain a greatly simplified structure, our construction coincides with Refs. 21, 22, 23. Our work has immediate relevance to recent cavity QED experiments involving two or more spin ensembles [24, 25, 26, 27].
@@
While the space of states accessible with our approach is vast, we focus on two particularly intriguing classes. In Secs. III and IV, we demonstrate how multi-ensemble entangled states with exceptional metrological properties for field gradients and curvatures can be naturally stabilized. Our approach solves an important challenge in differential quantum sensing [28, 16] by providing the first example of Heisenberg-limited differential sensing immune to common-mode noise and accessible with only simple one-body measurements - the standard measurement primitive in essentially all atomic sensors. In Sec. V, we discuss how our approach can stabilize states of fundamental interest in many-body physics: an entire family of entangled states in a 1D chain of spin-$S$ ensembles exhibiting symmetry-protected topological (SPT) order [29, 30] and connecting directly to recent ideas for exploiting sequential unitary circuits [31]. A special case of our protocol efficiently stabilizes the celebrated Affleck–Kennedy–Lieb–Tasaki (AKLT) state [32, 33]. Together, these examples highlight both the power and versatility of our general approach.
@@ II Setup for reconfigurable reservoir engineering
Our approach is sketched in Fig. 1. We start with a simple, fully-collective system where $N_{\rm tot}$ two-level atoms experience a single collective decay process (induced by resonant interaction with a lossy cavity), and are also subject to uniform and resonant Rabi drives. This realizes the well-known cooperative Rabi fluorescence (CRF) model [5, 6]. Working in the rotating frame of Rabi drives and adiabatically eliminating the cavity, the spin dynamics follows a Gorini–Kossakowski–Sudarshan–Lindblad (Lindblad) master equation:
@@
The CRF model is fully collective (i.e., total angular momentum is conserved), severely limiting the complexity of the resulting dynamics and steady states. As promised, we will achieve a richer dynamics by perturbing the system Hamiltonian, while keeping the dissipative part of the dynamics unchanged (see Fig. 1). We first partition our full set of $N_{\rm tot}$ atoms equally into $L$ sub-ensembles (indexed by $l=1,\dots,L$, with $L$ an even number). Then we break the permutation symmetry by assigning detuning $\delta_{l}$ for atoms in sub-ensemble $l$, and introducing an all-to-all spin-exchange interaction (amplitude $\chi$) with a “chiral” structure [21, 22, 23]. The modified master equation describes a system of $L$ ensembles with the same collective spin $S=N_{\rm tot}/2L$, and has the form:
@@
All the new terms introduced in $\hat{H}$ are compatible with the experimental tools of cavity QED. One can implement Eq. (2) by trapping many spin ensembles in a cavity as shown in Fig. 1. The detunings $\delta_{l}$ can be realized by applying separate Rabi drives with different frequencies that selectively address individual sub-ensembles, or by using a single global drive while engineering effective detunings for each ensemble via local frequency shifts (e.g. AC Stark shifts, Zeeman shifts). Similar to the setup in Ref. 24, the chiral spin-exchange couplings $\chi$ can be engineered by applying a magnetic-field gradient along the cavity, which allows for engineering cavity-assisted Raman processes between ensembles separated by different distances independently via frequency selection. Alternatively, one can also engineer $\chi$ using a chiral quantum network [21, 22, 23, 39, 40], which functions by coupling each spin ensemble to an optical cavity, and then connecting all these cavities via chiral waveguides. We discuss the details of the implementation of chiral spin-exchange couplings $\chi$ in App. G.
@@
Note that our approach is connected to (but distinct from) previous work exploiting non-reciprocal interactions for entanglement stabilization. In the limit where the chiral interaction is tuned to exactly $\chi=\Gamma$, Eq. (2) has the form of a cascaded master equation [41, 42, 21], and mimics a setup where the $L$ ensembles are all coupled to a common unidirectional waveguide (meaning that ensemble $l$ is only influenced by ensembles $k<l$). In the case $L=2$, dissipative pure-state entanglement in this fully directional limit can be understood as an example of the coherent quantum absorber method [21] and a consequence of hidden time-reversal symmetry [43]; applications to two spin ensembles was studied in Ref. 36, 44. Our work shows that for $L=2$, entanglement generation survives beyond the perfectly directional limit, indicating a kind of adiabatic continuity of the steady state as one tunes $\chi$.
In the other extreme case $L=N_{\rm tot}$ (i.e., each ensemble is just a single qubit with $S=1/2$), our work reduces to the entanglement stabilization scheme of Refs. 21, 22, 23. Note that, in this case, the complexity of the stabilized states is reduced, and there is no obvious metrological utility to the states that are generated. Although the possibility of states with SPT order exists for the case of $S=1/2$, it was not discussed by the previous works. Our work non-trivially generalizes this scheme to the case where each subsystem has an arbitrary spin $S>1/2$.
@@ III.2 Quantum enhancement for differential sensing
Figure 3: (a) Generalized Bloch sphere representation of the $L=2$ steady state. It can be interpreted as a two-mode spin squeezed state with squeezing along the $\hat{S}^{y}_{1}+\hat{S}^{y}_{2}$ and $\hat{S}^{z}_{1}+\hat{S}^{z}_{2}$ axes, and anti-squeezing along the $\hat{S}^{z}_{1}-\hat{S}^{z}_{2}$ and $\hat{S}^{y}_{1}-\hat{S}^{y}_{2}$ axes. Applying a differential phase $\phi$ between the two ensembles is equivalent to rotations about the $\hat{S}^{z}_{1}-\hat{S}^{z}_{2}$ axis in the first Bloch sphere. (b) Comparison between the optimal time scale to reach the steady state ($t_{\rm ss}$) and the steady-state squeezing (described by $\mathrm{Var}(\hat{S}^{y}_{1}+\hat{S}^{y}_{2})_{\rm ss}$), with $S=30$. Note that $\mathrm{Var}(\hat{S}^{y}_{1}+\hat{S}^{y}_{2})_{\rm SQL}=S$. We consider the initial state where all the spins are pointing down, and define $t_{\rm ss}$ as the optimal evolution time for the infidelity to the steady state to reach $10^{-3}$ with fixed $\Gamma$ and $\Delta/\Omega$. The same initial state is used in (c). (c) The scaling of optimal two-mode Wineland spin-squeezing parameter $\xi^{2}_{\rm opt}$ with spin $S$ for each ensemble and single-atom cooperativity $C$ (obtained from a second-order cumulant expansion). The scaling $\xi^{2}_{\rm opt}\propto 1/\sqrt{SC}$ is determined via numerical fitting.
@@
A central question is how this QFI scales with the number of atoms, i.e., with $S$. Fig. 2(d) shows this scaling for different choices of the parameter $\Delta/\Omega$. As discussed in the previous section, away from the large drive regime, i.e., when $\Delta/\Omega\gg 1/S$, the steady state’s angular-momentum distribution is localized near $J\sim S\Delta/\Omega$. It then follows from Eq. (10) and the form of the $B_{J,-J}$ coefficients that $\mathcal{F}_{\phi}\propto S\Omega/\Delta$. This immediately suggests how to achieve different canonical scalings of the QFI. For $\Delta/\Omega\sim 1$, we have a scaling near the standard quantum limit (SQL), $\mathcal{F}_{\phi}\propto S$. For a Rabi drive that increases with system size, $\Delta/\Omega\sim 1/\sqrt{S}$, we obtain a scaling that exceeds the SQL, $\mathcal{F}_{\phi}\propto S^{3/2}$.
For even larger Rabi drives, $\Delta/\Omega\sim 1/S$, we obtain a QFI that exhibits Heisenberg-limited (HL) scaling, i.e., $\mathcal{F}_{\phi}\propto S^{2}$. This is consistent with our earlier observation that our state approaches the $|J=0,m=0\rangle$ state in the large-Rabi-drive limit ($\Omega\rightarrow\infty$), a state which has a maximal QFI $\mathcal{F}_{\phi}]_{\rm max}=16S(S+1)/3$. The same maximal QFI is also reported in Refs. 15, 16. Since the $|J=0,m=0\rangle$ state is invariant under global rotations, in the large-Rabi-drive limit, one can obtain the same HL-scaling QFI for parameter estimation corresponding to generators $\hat{S}^{x}_{1}-\hat{S}^{x}_{2}$, $\hat{S}^{y}_{1}-\hat{S}^{y}_{2}$, and $\hat{S}^{z}_{1}-\hat{S}^{z}_{2}$, and vanishing variances for the operators $\hat{S}^{x}_{1}+\hat{S}^{x}_{2}$, $\hat{S}^{y}_{1}+\hat{S}^{y}_{2}$, and $\hat{S}^{z}_{1}+\hat{S}^{z}_{2}$. The squeezing and anti-squeezing of these operators can be extended to the case of a finite $\Omega$ as we show later, although the rotation symmetry of the three directions no longer holds.
@@
where $B_{J,m}^{-}=\langle J+1,m-1|\hat{S}^{-}_{1}-\hat{S}^{-}_{2}|J,m\rangle$. Since $\mathcal{F}_{\phi}\propto S\Omega/\Delta$, as discussed previously, we have $\langle\hat{S}^{x}_{1}-\hat{S}^{x}_{2}\rangle_{\rm ss}\propto S$ for a wide range of $\Delta/\Omega$, with corrections when decreasing $\Delta/\Omega$ below $\sim 1/S$. This result suggests the possibility of having the QFI exhibit Heisenberg-scaling while still having a non-zero average spin length of order $S$, e.g., by having $\Delta/\Omega\sim 1/S$. We indeed find that this is true through an explicit evaluation of Eq. (11). Note that having a spin length of order $S$ not only ensures that a Ramsey-type protocol is applicable, but also enhances the robustness against detection noise via large measurable signals.
Based on the large spin length $\langle\hat{S}^{x}_{1}-\hat{S}^{x}_{2}\rangle_{\rm ss}\propto S$ in our case, as shown in Fig. 3(a), one can further interpret our steady state $|\psi_{\rm ss}^{(2)}\rangle$ as a two-mode spin squeezed state (see Refs. 46, 47, 15 for other types of two-mode spin squeezed states). The squeezed quadratures are along the $\hat{S}^{y}_{1}+\hat{S}^{y}_{2}$ and $\hat{S}^{z}_{1}+\hat{S}^{z}_{2}$ axes, and the anti-squeezed quadratures are along the $\hat{S}^{z}_{1}-\hat{S}^{z}_{2}$ and $\hat{S}^{y}_{1}-\hat{S}^{y}_{2}$ axes. Both of the Bloch spheres share the same spin length, which is along the $\hat{S}^{x}_{1}-\hat{S}^{x}_{2}$ axis. In particular, the differential phase $\phi$ leads to rotations along the $\hat{S}^{z}_{1}-\hat{S}^{z}_{2}$ axis, resulting in a non-zero average value of $\hat{S}^{y}_{1}+\hat{S}^{y}_{2}$. This suggests a simple Ramsey-style measurement to measure the average collective spin component $\hat{S}^{y}_{1}+\hat{S}^{y}_{2}$. The quantum enhancement of this Ramsey-type measurement can be characterized by the two-mode generalization of the Wineland spin squeezing parameter [15]. Letting $(\Delta\phi)_{\rm est}$ denote the error in estimating $\phi$ from the Ramsey measurement, we have
@@
Note that $\mathrm{Var}(\hat{S}^{y}_{1}+\hat{S}^{y}_{2})_{\rm ss}=\sum_{J=0}^{2S}|c_{J,-J}|^{2}J/2$. Based on the previous discussions of $|c_{J,-J}|^{2}$, one can obtain $\xi_{\rm ss}^{2}\propto\Delta/\Omega$ when $\Delta/\Omega\gg 1/S$. Thus, by setting $\Delta/\Omega\sim 1/S$, one finds that Eq. (12) is approaching HL scaling, $\xi_{\rm ss}^{2}\propto 1/S$. We thus have established a major advantage of our scheme: HL scaling can be achieved for the differential phase $\phi$ using a simple Ramsey-type measurement with a signal of size $O(S)$. Based on the Bloch sphere representation, one can also extract the phase imprinted by $\hat{S}^{y}_{1}-\hat{S}^{y}_{2}$ with HL scaling by measuring $\hat{S}^{z}_{1}+\hat{S}^{z}_{2}$.
We stress that the quantum enhancement for differential sensing does not require exactly the same atom number for the two spin ensembles (see App. H for detailed analysis). For constant population imbalance, the HL scaling $\xi_{\rm ss}^{2}\propto 1/S$ is maintained. Even with $O(\sqrt{S})$ population imbalance, one can still obtain steady state spin squeezing $\xi_{\rm ss}^{2}\propto 1/\sqrt{S}$.
Our approach and the resulting steady state offer several distinct advantages over previous dissipative preparation schemes for enhanced differential sensing based on entanglement between two spin ensembles. Ref. 16 stabilizes a state with vanishing spin length, precluding its use in Ramsey-type measurements. Ref. 15 stabilizes a similar (but distinct) two-mode squeezed state, while it relies on engineering two independent dissipators, making the setup more demanding experimentally. In contrast, our scheme is naturally compatible with a simple Ramsey-type protocol and achieves HL scaling with a simpler experimental design.